HUBBLE SPACE TELESCOPE DAILY REPORT #2508 PERIOD COVERED: 0000Z (UTC) 11/24/99 - 0000Z (UTC) 11/29/99 Daily Status Report as of 333/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: Since HST remains in Zero Gyro Sun Point (ZGSP), no science observations were performed. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: No activity 2.2 FHST Updates: No activity 2.3 Operations Notes: The HST remains in ZGSP. The Canberra and Goldstone Deep Space Network sites were scheduled throughout the holiday weekend to dump engineering data to the ground for analysis. The SSR pointers were set twice, per ROP SR-3, in support of this task. CCS string "A" was recycled and operations were transferred to CCS string "G" at 328/2200Z. At 329/0136Z, operations returned to CCS string "A". Then, at 330/1233Z, the CCS string "A" FEP locked up (no real-time telemetry processed). Operations transferred to CCS "B" string at 330/1325Z and returned to "A" string at 330/1626Z. CCS string "B" was configured for GSTDN operations at 330/1807Z, although no GSTDN support was utilized on day 330 due to CCS string problems. CCS string "B" was re-booted at 332/1850Z. Using ROP SR-1, the SSR EDAC error counter was cleared four times. The "Periodically Reset Time-Based Recharge Ratio Charge/Discharge Buffer" was adjusted five times in accordance with Operations Note #661. GRG1_2MC broke the limit setting at 329/044922Z and at 329/045322Z. HSTAR 7380 was written. Operations Note #674 was used to "Adjust the Limit for GRG1_2MC" at 329/0552Z. Per HSTAR 7381, the charge rate error counter (QROCERC) flagged out-of-limits high with a value of 579 at 330/064721Z, returning within boundaries (count of 500) at 330/064920Z. The parameter flagged again at 330/082421Z with a value of 536. It returned to nominal ranges at 330/081619Z. Four NICMOS mnemonics were out-of-limits low by one or two counts in the first hour of day 331. The parameters were: NBSHLTMP, NFOBTMP, NFTRSTMP, and NTSHLTMP. HSTAR 7382 was written. Operations Note #675 was utilized at 332/2218Z to perform "SI Limit Management". A full DF-224 memory dump was performed at 333/0636Z as directed by ROP DF-1. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Servicing Mission 3A, currently scheduled for December 9, 1999. /CAW